OpenStudy (anonymous):

What is the simplified form of the square root of the quantity 72 times n to the thirteenth power.? 8nthe square root of the quantity 3 times n to the sixth power. 3n6the square root of the quantity 8 times n. 8n6the square root of the quantity 3 times n. 6n6the square root of the quantity 2 times n.

OpenStudy (anonymous):

First of all you take \[\sqrt{72}\] then you will get \[6\sqrt{2}\] and then you take \[\sqrt{n^{13}}\] which will be\[n^6\sqrt{n}\] cus \[\sqrt{n^{13}} = n^{\frac{ 13 }{ 2 }} \] which is \[n^6\sqrt{n}\] then the answer will be the last chioce (D) Let me know if you got it :)

OpenStudy (anonymous):

thank you, i got it!